in DNA of E coli has 20% bases of guanine. what will be fraction of thymine
  1. 30
  2. 3
  3. 15
  4. 5

Grade:12

1 Answers

Gaurav
9 years ago

Hello student
Correct answer is 1. 30

G = 20%
As, C=G
C= 20%
C+G = 20+20 = 40%
We Know,
A+T = 100-40 = 60
Also,
A=T ⇒ T = 60/2 = 30 ans....